1. Formaldehyde can be produced by oxidation of C2H4 and the following reaction may occur. C2H4 + ½ O2  C2H4O C2H4 + 3O2  2CO2 + 2H2O The product gas analysis shows 28.57% C2H4O, 11.43% CO2, 11.43% H2O, 22.86% C2H4 and 25.71 % O2. Calculate the % conversion of C2H4 and selectivity, C2H4O / CO2. Use extent of reaction method.
Answer #1

Ca Hut / a Sol Chemical reachon, Chuo Cyty + 30, 200, +240 the product gas analysis, 28.57% Gry 11.03 o lo Co 11.43 %. Mao 22Hached 244 = 28.57 +5.115 = 340285 mol 22.86 mol Unreached GHY 34.285+ 22.106 = 57.145 mal total Qty feel 9% conversion of an